# Image Demoireing in RAW and sRGB Domains

### 14 Dec 2023 v1

@@ 3.1 Design of SCDM
Params (M) 1.55 3.36 5.93 6.04 2.80 5.33 2.68 2.46
TFLOPs 0.095 0.882 4.478 0.161 0.013 0.341 0.221 0.093
Inference time 0.052 0.284 0.115 1.094 0.070 0.182 0.058 0.089
FSM. For sRGB demoiréing, we leverage a learnable band stop filter [32] to mitigate moiré patterns in the frequency domain. Considering the formation of moiré patterns is closely related to the camera imaging process, especially the frequency of the color filter array, demoiréing in the frequency domain becomes a favorable approach. However, the frequency spectrum of the moiré pattern is often entangled with that of the original contents, thereby increasing the difficulty of disentanglement. Therefore, we propose to perform the disentanglement within small patches, where the moiré patterns tends to be more pronounced in a specific range of frequency bands, thereby facilitating the removal of moiré patterns. Inspired by Zheng et al. [32], we utilize a band stop filter to amplify certain frequencies and suppress others using Block-DCT. However, obtaining the frequency spectrum prior to model the moiré texture is challenging, and determining a frequency prior for each moiré image patch is time-consuming. To address this issue, a set of convolution layers are employed as the learnable band stop filter $\mathrm{B}(\cdot)$ to attenuate the specific frequencies of moiré patterns while preserving the original image contents. The process of the FSM can be represented as:
where $\mathrm{DCT}$ and $\mathrm{IDCT}$ denote the process of Block-DCT and the corresponding inverse process.
3.2 Design of RGISP
RGISP is proposed to transform pre-demoiréd RAW features $\mathbf{D}_{raw}$ into sRGB domain $\mathbf{D}_{out}$, with the assistance of pre-demoiréd sRGB features $\mathbf{D}_{rgb}$. Matrix transformation is commonly employed in the traditional ISP pipelines [1].It enables the enhancement or conversion of image colors to another color space through a channel-wise matrix transformation, which is facilitated by globally shared settings like environmental illumination and color space specifications [9]. Building upon this principle, we introduce matrix transformation as a means to perform color transformation, as illustrated in Fig. 3. The design of this matrix transformation is informed by recent advancements in transposed cross-attention.
Figure 3: The structure of RGISP.
Given $\mathbf{D}_{raw}\in\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}$, the vectors of the query $\mathbf{Q}\in\mathbb{R}^{C\times\frac{H}{2}\frac{W}{2}}$, key $\mathbf{K}\in\mathbb{R}^{\frac{H}{2}\frac{W}{2}\times C}$ are generated through the projection with a 1 $\times$ 1 convolutional layer $\mathrm{Conv}$ and a flatten operation $\mathrm{Flatten}$, where $C$ is the intermediate channel number. Similarily, value $\mathbf{V}\in\mathbb{R}^{C\times\frac{H}{2}\frac{W}{2}}$ can be acquired by $\mathbf{D}_{rgb}\in\mathbb{R}^{C\times\frac{H}{2}\times\frac{W}{2}}$ as:
$\mathbf{Q},\mathbf{K}=\mathrm{Flattern}(\mathrm{Conv}(\mathbf{D}_{raw})),$ (4)
$\mathbf{V}=\mathrm{Flatten}(\mathrm{Conv}(\mathbf{D}_{rgb})).$ (5)
where a scaling coefficient $\lambda$ is applied for numerical stability. Subsequently, the vector $\mathbf{V}$ is transformed by the matrix $\mathbf{M}$, performing color space conversion in feature-level. The output feature after color transformation can be obtained by $\mathbf{D}_{out}=\mathbf{M}\cdot\mathbf{V}$. As a complement to the global matrix transformation, we use a depth-wise convolution and a point-wise convolution to refine the local details.
3.3 Loss Function
To effectively remove the moiré patterns and adjust the color deviations, we introduce supervision on different domains, clean RAW and clean sRGB. The overall training objective can be expressed as:
where ${\mathbf{Y}}_{raw}$ and ${\mathbf{Y}}_{rgb}$ are the ground truth RAW image and sRGB image. Only $L_{1}$ loss is employed for both RAW and sRGB supervision. Since our goal is to restore the image in the sRGB domain, we empirically set $\alpha$=0.5.
4 Experiments
4.1 Experimental Setup
Dataset. The experiments are conducted on the RAW image demoiréing dataset published in TMM22 [27], which contains recaptured scenes of natural images, documents, and webpages. The moiré images are captured using four different smartphone cameras and three display screens of varying sizes. In TMM22, the dataset consists of 540 RAW and sRGB image pairs with ground-truth for training, and 408 pairs for testing. To facilitate the training and comparison process, patches of size 256$\times$256 and 512$\times$512 are cropped respectively for the training and testing sets.
Training Details. The proposed RRID utilizes both RAW and sRGB images as inputs to generate the restored image in the sRGB domain. The model is trained using the AdamW optimizer with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. A multistep learning rate schedule is employed, with the learning rate initialized at 2×$\times 10^{-4}$. The RRID model is trained for 500 epochs using a batch size of 80 on 4 NVIDIA RTX 3090 GPUs.
4.2 Quantitative Results
We compare our approach with several image demoiréing methods in sRGB domain, including DMCNN [20], WDNet [15], ESDNet [26], a RAW image demoiréing method RDNet [27]. We further add comparison with a low-light enhancement method DNF [9] with RAW input, a video demoiréing method RawVDmoiré [29] with RAW input, and a deflection removal method CR3Net [19] with paired RAW and sRGB data. To evaluate the performance of our proposed RRID, we adopt the following three standard metrics to assess pixel-wise accuracy and the perceptual quality: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [23], and Learned Perceptual Image Patch Similarity (LPIPS) [30]. Additionally, the number of parameters, Floating-Point Operations per Second (FLOPs), and inference time are introduced to compare the model complexity. To ensure a fair comparison, we adjust the inputs of certain models to align with the dataset requirements. Subsequently, all models are fine-tuned using the default parameter settings as provided in their paper on the TMM22 dataset. The superior results are reported between the pre-trained and fine-tuned models.
@@ 4.2 Quantitative Results
Table. 1 presents the quantitative and the computational complexity comparisons. Our proposed RRID outperforms the second-best method, RawVDmoiré, 0.62dB in PSNR and 0.003 in SSIM. As for LPIPS, our methods accomplishes an acceptable result at 0.079, surpassing most of the comparison methods. Even though our method utilizes both RAW data and sRGB data as inputs, our approach can still achieve a satisfactory demoiréing effect with a inference time of merely 0.089s. Compared to previous methods, our RRID demonstrates superiority by employing relatively fewer parameters and FLOPs, further demonstrating the superiority of our approach.
@@ 4.3 Qualitative Results
We present visual comparisons between our approach and the existing methods in Fig. 4. The results demonstrate the advantage of RRID in effectively removing moiré artifacts and correcting color deviations. For the first scene demonstrating a recaptured flower with moiré patterns, previous methods tend to exhibit residual moiré patterns in their restored images. While ESDNet successfully eliminates a significant portion of these patterns, it struggles to accurately restore the original colors of the image. Apart from natural scenes, we also present a case of webpages contaminated with moiré patterns. Through the utilization of both RAW and sRGB data, our approach adeptly eliminates moiré patterns while preserving fine image details, even in scenarios characterized by severe color deviations. This ensures a more accurate restoration of the original colors and results in enhanced visual effects. Additional qualitative comparisons can be found in the supplementary file.
@@ 4.4 Ablation Study
1) Ablation study on the model architecture and inputs.
Table 3: Baseline models comparison for the evaluation of the architecture and the inputs.
[…]
$B_{1}$ (w/o RAW in and RAW branch) 25.79 0.915
$B_{2}$ (w/o sRGB in and sRGB branch) 27.24 0.929
$B_{3}$ (w/o RAW in) 26.51 0.923
$B_{4}$ (w/o RGISP) 27.38 0.932
$B_{5}$ (w/o Reconstruction) 26.93 0.926
$B_{6}$ (full version) 27.88 0.938
In Table 3, we investigate the architecture and inputs of RRID and validate the importance of different individual components in the whole RRID. We conduct this ablation study by comparing the proposed RRID and the following variants of RRID:
@@
    $B_{1}$. Both the RAW input and the RAW demoiréing branch are removed.
@@
    $B_{3}$. The sRGB input is substituted with RAW input. The input channel number is adjusted, while the rest of the architecture is unchanged.
@@
    $B_{6}$. The full model of RRID with both RAW and sRGB images as model input.
@@
Our proposed RRID leverages both RAW data and sRGB as input, employing distinct demoiréing modules for each data type to extract pre-demoiréd features. Removing the RAW input and its corresponding raw demoiréing module ($B_{1}$) results in a drastic decrease in PSNR of approximately 2dB. Likewise, eliminating the sRGB input and its corresponding RGB demoiréing module ($B_{2}$) leads to a PSNR of only 27.24dB. In addition to the decline in numerical metrics, we can also assess the visual performance from Fig. 5. When relying only on sRGB data and its corresponding demoiréing modules, moiré patterns are less effectively removed, as depicted in Fig.5(b). Conversely, Fig.5(c) demonstrates that using only raw data as input results in images with more pronounced color deviations. To further validate the significance of RAW input for image demoiréing, we compare ($B_{3}$) to ($B_{1}$) by solely removing the RAW input data while retaining the remaining model structures. Although ($B_{3}$) demonstrates some improvement in numerical metrics compared to ($B_{1}$), its performance remains unsatisfactory due to the lack of the additional rich original information provided by the RAW input and the fact that the original raw demoiréing module is specifically designed for RAW data. Additionally, we conducted further comparisons by removing the originally designed RGISP ($B_{4}$) and Reconstruction modules ($B_{5}$), resulting in inferior numerical metrics and performance in Fig. 5 compared to the complete RRID model ($B_{6}$).
2) Ablation study on SCDM.
@@
SCDM is designed to perform efficient demoiréing on both raw data and sRGB data, with distinct modules incorporated into skip-connections.More specifically, for raw feature demoiréing, we introduce GFM into the skip connection, enabling adaptive differentiation between texture details and moiré patterns through feature gating. Similarily, FSM is embeded in SCDM for sRGB moiréd features, leveraging a learnable band stop filter to mitigate moiré patterns in the frequency domain.
@@
Table 4 demonstrates the effectiveness and indispensability of the meticulously designed GFM and FSM integrated within the skip-connections. Removing the demoiréing modules on the skip-connections ($S_{1}$-$S_{3}$) individually results in varying degrees of performance decline for the model. Fig.6(b) demonstrates that the model’s ability to remove moiré patterns decreases upon removing the GFM and FSM modules. Furthermore, to assess the specificity of GFM and FSM, we substitute the original modules in $S_{4}$ and $S_{5}$, respectively. Despite no reduction in the number of parameters, the PSNR decreased by approximately 0.3-0.5dB. Since DCAB is a crucial component module in SCDM, we substitute DCAB with two convolutional layers to confirm its necessity. Fig. 6(c) showcases more residual moiré patterns and significant color deviation compared to our complete SCDM presented in Fig.6(d).
3) Ablation study on RGISP.
@@
RGISP is designed to integrate color information from the coarsely demoiréd sRGB features during the ISP stage, facilitating the color recovery process. The process of the RGISP can be recognized as a cross-attention mechanism. To assess the efficacy of the RGISP design, we compare the following structures in Table 5. For $R_{1}$, we concatenate the pre-demoiréd RAW features with the sRGB features and subsequently apply a self-attention mechanism. In $R_{2}$, the pre-demoiréd RAW features undergo a self-attention mechanism and are then concatenated with the pre-demoiréd sRGB features. Table 5 shows that our proposed RGISP surpasses the performance of self-attention mechanisms ($R_{1}$, $R_{2}$), as well as other raw-to-rgb conversion methods RRM [19] ($R_{3}$), based on numerical metrics. Fig. 7 depicts a webpage with globally distributed moiré patterns. $R_{1}$-$R_{4}$ are all capable of effectively removing the moiré patterns, but $R_{4}$ exhibits the best performance in correcting color deviation. With the introduction of sRGB pre-demoiréd features, RGISP is able to restore the original colors of the image effectively.
@@ 5 Conclusion
In this paper, we propose to concurrently utilize both RAW and sRGB data for image demoireing (RRID). Firstly, we introduce SCDM with dedicated modules embedded in skip-connections to handle the demoireing of RAW and sRGB features, respectively. Additionally, we present RGISP, which integrates color information from coarsely demoired sRGB features into the ISP stage to aid in the color recovery process. Extensive experiments demonstrate the superior performance of our RRID method in both moire pattern removal and color cast correction. Our approach outperforms state-of-the-art methods by 0.62dB in PSNR and 0.003 in SSIM.
@@ 6 Additional Ablation Study for SCDM
In Table 6, we validate the effectiveness of embedding the demoiréing modules in the skip-connections. We conduct this ablation study by comparing the proposed SCDM and the following cases:
@@
    $A_{1}$. In the Raw branch, DCAB is replaced with GFM, while in the RGB branch, DCAB is replaced with FSM. The demoiréing modules embedded in the skip-connections are reserved for both branches.
@@
    $A_{2}$. In the Raw branch, DCAB is replaced with GFM, while in the RGB branch, DCAB is replaced with FSM. The demoiréing modules embedded in skip-connection are removed for both branches.
@@
From $A_{1}$, it can be observed that replacing DCAB with GFM and FSM results in a noticeable decline in both PSNR and SSIM. A more significant concern arises from the substantial increase in inference time, primarily caused by the block-DCT utilized in FSM. This observation underscores the crucial role of DCAB, a pivotal block in SCDM, in encoding and decoding informative features. Moreover, even when we eliminate the demoiréing modules on the skip-connections in $A_{1}$, the inference time remains relatively long for $A_{2}$. This finding emphasizes the effectiveness and efficiency of incorporating demoiréing modules within skip-connections.
@@ 7 Limitation
While we propose employing both RAW and sRGB images for moiré pattern removal, and design RGISP to integrate color information from coarsely demoiréd sRGB features with demoiréd RAW features during the ISP stage, achieving flawless color restoration in scenes with significant color deviations remains a challenging task.
@@
Fig. 8 demonstrates two scenes where the input images containing moiré patterns display noticeable color discrepancies compared to the ground truth images. While our proposed method effectively removes moiré patterns and partially corrects the color, it falls short of fully mapping the colors to those present in the ground truth. Currently, there is only one dataset, TMM22 dataset [27], available for image demoiréing with RAW-RGB pairs. However, the images in the training set of TMM22 are limited to a size of 256$\times$256, which may constrain the model’s capacity to learn global color mapping. To overcome this limitation, one possible solution is to collect a RAW-RGB paired image demoiréing dataset with a higher resolution, which will be our future work.
@@ 8 More Qualitative Comparison on TMM22 Dataset
More visual comparisons with state-of-the-arts methods are presented in Figs. 9-10. The results demonstrate the superiority of our proposed RRID in demoiréing and restoring color. For instance, in the first scene depicted in Fig. 9, the input image suffers from both severe moiré pattern contamination and noticeable color deviation. Previous methods struggle to remove moiré patterns and encounter difficulties in accurately restoring the petal color to pink. In contrast, our approach successfully eliminates moiré patterns and effectively corrects color deviations. The last scene illustrated in Fig. 10 presents a moiréd document. Given the relatively slight moiré contamination in this scene, the majority of methods successfully remove the moiré patterns. Notably, the colored area at the top of the image deserves attention as most methods incorrectly correct it as blue. However, the ground truth image reveals that this area is actually cyan. In contrast, our method excels at faithfully restoring the image’s color.

### 15 Mar 2024 v2

@@ 4.2 Quantitative Results
Params (M) 1.55 3.36 5.93 6.04 1.25 5.33 2.68 2.38
TFLOPs 0.102 0.055 0.141 0.161 0.013 0.022 0.883 0.093
Inference time (s) 0.052 0.284 0.115 1.094 0.070 0.182 0.058 0.089
@@
Table. 1 presents the quantitative and the computational complexity comparisons. RRID outperforms the second-best method, RawVDmoiré, by 0.62dB in PSNR and 0.003 in SSIM. As for LPIPS, our methods accomplishes the second-best result at 0.079, surpassing most of the comparison methods. Even though RRID utilizes both RAW and sRGB data as inputs, our approach can still achieve a satisfactory demoiréing effect with a inference time of merely 0.089s. Compared to previous methods, our RRID demonstrates superiority with acceptable parameters and FLOPs, further demonstrating the superiority of our approach.
@@ 4.3 Qualitative Results
We present visual comparisons among our approach and the existing methods in Fig. 4. The results demonstrate the advantage of RRID in effectively removing moiré artifacts and correcting color deviations. For the first scene demonstrating a recaptured flower with moiré patterns, previous methods tend to exhibit residual moiré patterns in their restored images. While ESDNet successfully eliminates a significant portion of these patterns, it struggles to accurately restore the original colors of the image. Apart from natural scenes, we also present a case of webpages contaminated with moiré patterns. Through the utilization of both RAW and sRGB data, our approach eliminates moiré patterns while preserving fine image details, even in scenarios characterized by severe color deviations. This ensures a more accurate restoration of the original colors and results in enhanced visual effects. Additional qualitative comparisons can be found in the supplementary file.
@@ 4.4 Ablation Study
1) Ablation study on the model architecture and inputs. In Table 3, we investigate the architecture and inputs of RRID and validate the importance of different individual components in the whole RRID. We conduct this ablation study by comparing the proposed RRID and the following variants of RRID:
@@
    $B_{1}$. Both the RAW input and the RAW demoiréing branch are removed. Due to the lack of RAW features, the entire RGISP is also deleted.
@@
    $B_{3}$. The sRGB input is substituted with RAW input. The input channel number is adjusted, while the rest of the architecture is kept unchanged.
@@
    $B_{6}$. The full model of RRID with both RAW and sRGB images as input.
Table 3: Comparison of baseline models for the evaluation of architecture and inputs.
[…]
$B_{1}$ (w/o RAW in and RAW branch) 25.79 0.915
$B_{2}$ (w/o sRGB in and sRGB branch) 27.24 0.929
$B_{3}$ (w/o sRGB in) 26.51 0.923
$B_{4}$ (w/o RGISP) 27.38 0.932
$B_{5}$ (w/o Reconstruction) 26.93 0.926
$B_{6}$ (full version) 27.88 0.938
Our proposed RRID leverages paired RAW-sRGB data as input. Removing the RAW input and its corresponding RAW demoiréing module ($B_{1}$) results in a drastic decrease in PSNR of approximately 2dB. Likewise, eliminating the sRGB input and its corresponding RGB demoiréing module ($B_{2}$) leads to a PSNR of only 27.24dB.
@@
In addition to the decline in numerical metrics, we can also assess the visual performance from Fig. 5. When relying only on sRGB data and its corresponding demoiréing modules, moiré patterns are less effectively removed, as depicted in Fig.5(b). Conversely, Fig.5(c) demonstrates that using only RAW data as input generates images with more pronounced color deviations. To further validate the significance of RAW input for image demoiréing, we compare $B_{3}$ to $B_{1}$ by solely removing the RAW input data while retaining the remaining model structures. Although $B_{3}$ demonstrates some improvements in numerical metrics compared to $B_{1}$, its performance remains unsatisfactory due to the lack of information provided by the RAW input and the fact that the original RAW demoiréing module is specifically designed for RAW data. Additionally, we conducted further comparisons by removing the originally designed RGISP ($B_{4}$) and Reconstruction modules ($B_{5}$), resulting in inferior numerical metrics and performance in Fig. 5 compared to the complete RRID model ($B_{6}$).
2) Ablation study on SCDM. SCDM is designed to perform effective demoiréing on both RAW and sRGB features. More specifically, we introduce GFM for RAW features demoiréing and FSM for sRGB features demoiréing.
@@
Table 4 demonstrates the effectiveness and indispensability of the meticulously designed GFM and FSM integrated within the skip-connections. Removing the demoiréing modules on the skip-connections ($S_{1}$-$S_{3}$) individually result in varying degrees of performance decline of the model. For instance, removing both GFM and FSM leads to only 27.00 dB in PSNR. Furthermore, Fig.6(b) demonstrates that the model’s ability to remove moiré patterns decreases upon removing the GFM and FSM modules. Furthermore, to assess the specificity of GFM and FSM, we substitute the original modules in $S_{4}$ and $S_{5}$, respectively. Despite no reduction in the number of parameters, the PSNR decreases by approximately 0.3-0.5dB. Since DCAB is a crucial component module in SCDM, we substitute DCAB with two convolutional layers to confirm its necessity. Fig. 6(c) showcases more residual moiré patterns and significant color deviation compared to our complete SCDM presented in Fig.6(d).
@@
3) Ablation study on RGISP. RGISP is designed to integrate color information from the coarsely demoiréd sRGB features during the ISP stage, facilitating the color recovery process. The process of the RGISP can be recognized as a cross-attention mechanism. To assess the efficacy of the RGISP design, we compare the following structures in Table 5. For $R_{1}$, we concatenate the pre-demoiréd RAW features with the sRGB features and subsequently apply a self-attention mechanism. In $R_{2}$, the pre-demoiréd RAW features undergo a self-attention mechanism and are then concatenated with the pre-demoiréd sRGB features. Table 5 shows that our proposed RGISP surpasses the self-attention mechanisms ($R_{1}$, $R_{2}$), as well as other RAW-to-sRGB conversion methods RRM [19] ($R_{3}$).
@@
In addition, Fig. 7 depicts a webpage with globally distributed moiré patterns. $R_{1}$-$R_{4}$ are all capable of effectively removing the moiré patterns, but $R_{4}$ exhibits the best performance in correcting color deviation. With the introduction of sRGB pre-demoiréd features, RGISP is able to learn a device-dependent ISP and restore the original colors of the image effectively.
@@ 5 Conclusion
In this paper, we propose to jointly utilize RAW and sRGB data for image demoiréing (RRID). Firstly, we introduce SCDM with GFM and FSM embedded in skip-connections to handle the demoiréing of RAW and sRGB features, respectively. Additionally, we present RGISP to learn a device-dependent ISP, aiding the color recovery process. Extensive experiments demonstrate the superior performance of our RRID method in both moiré pattern removal and color cast correction. Our approach outperforms state-of-the-art methods by 0.62dB in PSNR and 0.003 in SSIM.
@@
In this supplementary material, we first conduct the additional ablation study for SCDM in Sec. 6. Then, a limitation of our proposed RRID is discussed in Sec. 7. Lastly, more visual comparisons between RRID and several state-of-the-art models are provided in Sec. 8.
@@ 6 Additional Ablation Study for SCDM
Table 6 presents the effectiveness and efficiency of the special design in SCDM, where the demoiréing modules, GFM and FSM, are embedded in the skip-connections of SCDM. In this ablation study, we compare the proposed SCDM with the following cases:
@@
    $A_{1}$. In the RAW branch, DCAB is replaced with GFM, while in the RGB branch, DCAB is replaced with FSM. The demoiréing modules embedded in the skip-connections remain unchanged for both branches.
@@
    $A_{2}$. In the RAW branch, DCAB is replaced with GFM, while in the RGB branch, DCAB is replaced with FSM. The demoiréing modules embedded in skip-connections are removed for both branches.
@@
From $A_{1}$ in Table 6, it can be observed that replacing DCAB with GFM and FSM results in a noticeable decline by 0.77dB in PSNR and 0.011 in SSIM. A more significant concern arises from the substantial increase in inference time to 4.589 seconds, primarily caused by the Block DCT utilized in FSM. This observation underscores the crucial role of DCAB in encoding and decoding informative features as a pivotal block in SCDM. Moreover, even when we remove the demoiréing modules from the skip-connections in $A_{1}$, the inference time remains relatively long for $A_{2}$. This observation emphasizes the effectiveness and efficiency of incorporating demoiréing modules within skip-connections.
@@ 7 Limitation
While we propose to utilize paired RAW and sRGB images for moiré pattern removal and design RGISP to learn a device-dependent ISP, performing color restoration in scenes with significant color deviations remains a challenging task.
@@
Fig. 8 demonstrates two scenes where the moiréd input images exhibit noticeable color discrepancies compared to the ground-truth images. While our proposed method effectively removes moiré patterns and partially corrects the color, it falls short of fully mapping the colors to those present in the ground-truth images. Currently, there is only one dataset, TMM22 dataset [28], available for image demoiréing with the RAW-sRGB image pairs. However, the images in the training set of TMM22 are limited to a size of 256$\times$256, which may constrain the model’s capacity to learn global color mapping. To overcome this limitation, one possible solution is to collect a new paired RAW-sRGB image demoiréing dataset with a higher resolution, which will be our future work.
@@ 8 More Qualitative Comparison on TMM22 Dataset
More visual comparisons with state-of-the-arts methods are presented in Figs. 9-10. The results demonstrate the superiority of our proposed RRID in demoiréing and restoring color. For instance, in the first scene depicted in Fig. 9, the input image suffers from both severe moiré pattern contamination and noticeable color deviation. Previous methods struggle to remove moiré patterns and encounter difficulties in accurately restoring the petal color to pink. In contrast, our approach successfully eliminates moiré patterns and restores the original color.
The last scene depicted in Fig. 10 presents a moiréd document. Despite the relatively subtle moiré contamination in this scene, the majority of methods successfully eliminate the moiré patterns. However, the colored area at the top of the image deserves attention, as most methods incorrectly correct it as blue. In contrast, the ground truth image reveals that this area is actually cyan. Remarkably, our method excels in accurately restoring the original color.